OpenStudy (ttp):

Micah filled his garden with 6 cubic feet of organic dirt. The garden is a rectangular prism with a volume V equals l times w times h where l is length, w is width, and h is depth. If the garden is 1 foot deep and 3 feet wide, how long is the garden? 1 foot 2 feet 3 feet 4 feet

OpenStudy (ttp):

@e.cociuba

OpenStudy (ttp):

Can you help me again please :) ? @e.cociuba

OpenStudy (anonymous):

2 feet because l x w x h = ? x 3 x 1 1 x 3 = 3 so if you want 6 feet, what is 6/3?

OpenStudy (e.cociuba):

V= l * w * h V = 6, h = 1 , w = 3 Plug the numbers into the formula: 6 = l * 3 * 1 l = 6/3 l = 2 feet